Horizon Value of Free Cash Flows

Current and projected free cash flows for Radell Global Operations are shown below.

Actual
2016

2017
Projected
2018

2019
Free cash flow $606.82 $667.50 $707.55 $750.00
(millions of dollars)

Growth is expected to be constant after 2018, and the weighted average cost of capital is 11.55%. What is the horizon (continuing) value at 2019 if growth from 2018 remains constant? Round your answer to the nearest dollar. Do not round intermediate calculations.
Answer #1

g = [FCF(2019) / FCF(2018)] - 1

= [750 / 707.55] - 1 = 1.06 - 1 = 0.06, or 6%

HV(2019) = [FCF(2019) * (1 + g)] / [WACC - g]

= [750 * (1 + 0.06)] / [0.1155 - 0.06] = 795 / 0.0555 = $14,324.27, or $14,324